### v1

# Is a Dissipative System Always a Gradient System or a Gradient Like System?

Abstract
We find that to the dynamics of a given dissipative system a p=1 differential form can be associated with a general decomposition into a potential term and a non-potential residual part. If the residual part is absent the form is closed and the system is gradient system or gradient like. If it is non-closed, in the differential form approach, it remains non-closed under a variable change of coordinates, i.e., the system is not a gradient one or a gradient like in any coordinate system. On the other hand, there are claims that a potential should always exists, i.e., the class of dissipative systems and the the class of gradient systems should coincide. We fix this conundrum by introducing a generalized change of coordinates that aims a transformation to a gradient system or a gradient like system. The condition of being closed in the new coordinates of a certain, through the generalized change of coordinates defined differential form, results in a nonlinear differential equation together with a consistency condition. We give examples of physical systems where an analytical solution for the transformation can be found, and hitherto, the potential, but even when the potential is not accessible analytically, we find that it always exist, and therefore we give in principle an affirmative answer to the defining question of this work. Our findings removes loopholes in the question if a potential may exist but it is not known.
pacs:
74.81.Fa,05.10.Gg
I Introduction.
Dynamical systems represented by a vector $\vec{x}$ in $\in R^{N}$, and whose evolution is given by stochastic equation(SDE):
$\dot{\vec{x}}=\frac{d\vec{x}}{dt}=\vec{f}(\vec{x})+\vec{g}(\vec{x})\zeta(\vec{x},t)$ (1)
constitutes a universal matter of study [1]. Here each of the functions $f(\vec{x})$ and $g(\vec{x})$ has sufficient well behavior such that ODE to have solutions, and the function $\zeta(t)$ represent an stochastic process (noise)[2]taken to be a Markov process. This equation without the noise term, has been intensively studied for the nature of the dynamics in particular, when the system is dissipative[3], [4][5]. Another issue is concerned with the existence of Lyapunov functions and its role in the classification of the dynamics[7]. If there exist a scalar function $V(\vec{x})=V(x_{i})$, $i=1,..N$, which is usually called the potential, such that it is bounded from below, and stationary on the limits set of the dynamical systems [3] and such that:
$\dot{\vec{x_{i}}}=f_{i}(\vec{x})=-\frac{\partial V(\vec{x})}{\partial x_{i}}\Rightarrow\frac{dV(\vec{x})}{dt}=-\sum_{i=1}^{N}\frac{\partial V(\vec{x})}{\partial x_{i}}\frac{\vec{x}_{i}}{dt}=\sum_{i=1}^{N}\left(\frac{\partial V(\vec{x})}{\partial{x_{i}}}\right)^{2}\leq 0,$ (2)
the potential is called a Lyapunov function for the system [3]. One can consider a more general situation given in which in the governing dynamics a residual vector function $\vec{\it{v}}$} appears, in addition to a positive definite matrix $\mathcal{S}(\vec{x})$:
$\dot{\vec{x_{i}}}=\vec{f}_{i}(\vec{x}):=-\sum_{j=1}^{N}\mathcal{S}_{ij}\frac{\partial\it{V}(\vec{x})}{\partial x_{j}}+\vec{\it{v}}_{i}(\vec{x}),$
${\vec{\it{v}}}\cdot\nabla\vec{V}=0$ (3)
i.e., the vector functions $\vec{\it{v}}$} and $\vec{\nabla}V(\vec{x})$ are orthogonal in the configuration space $U$ $\subset R^{N}$ (orthogonality condition). In this case the potential $\it{V}(\vec{x})$ still behaves as a Laypunov function, i.e., $\frac{dV(\vec{x})}{dt}\leq 0$. The orthogonality condition can be written, by directing rewriting it in the form:
$\left(\vec{f}(\vec{x})+\mathcal{S}\vec{\nabla}V(\vec{x})\right)^{T}\cdot\vec{\nabla}V(\vec{x})=\vec{f}^{T}(\vec{x})\vec{\nabla}V(\vec{x})+(\vec{\nabla}V(\vec{x}))^{T}\mathcal{S}\vec{\nabla}V(\vec{x})=0$ (4)
Previous equation, giving the orthogonality condition and the function $\vec{f}(\vec{x})$, can be though as an equation for determining $V(\vec{x})$ [9]. This is Graham and Tel[6] approach, where the existence of a globally defined potential function with respect to the positive define matrix $\mathcal{S}$ is defined. The residual part $\vec{\it{v}}$}, does not participate in the relaxation dynamics. When considering Equ.(1) for the special case of a Gaussian random processes, i.e.:
$\dot{\vec{x}_{i}}=\vec{f}_{i}(\vec{x})+\sum_{j}^{N}\vec{g}_{ij}(\vec{x})\zeta_{j}(t);\hskip 28.45274pt<\zeta_{i}(t)\zeta_{j}(t)>=2\epsilon\delta(t-t^{\prime}),$
$\vec{f}_{i}=-\sum_{j=1}^{N}\mathcal{S}_{ij}\frac{\partial\it{V}(\vec{x})}{\partial x_{j}}+\vec{\it{v}}_{i}\hskip 184.9429pt$
The Ansatz for $\vec{f}_{i}$ in previous equation fulfills the orthogonality condition. Now, the question for finding the potential (see [8] for a review), when considering the associated Fokker-Planck to Equ.(5), and other assumptions, namely, the validity of the fluctuation dissipation relation for the case when $\mathcal{S}$ is a constant matrix, when $\mathcal{S}=gg^{T}$, and the validly of the orthogonal relation and the divergenz free of $\vec{v}$, then there exist a stationary distribution $P_{stat}$ for the associated Fokker-Plank-equation given by in the small noise limit (lim $\epsilon\rightarrow 0$):
$P_{stat}=Nexp\left(-\frac{V(\vec{x})}{\epsilon}\right)$ (6)
from which one could obtain the potential if $P_{stat}$ could be independently obtained, as the limit:
$V(\vec{x})=lim_{\epsilon\rightarrow 0}(-\epsilon\ln{P_{stat}})$ (7)
also known as Graham potential. Graham and Tel [6] have pioneered the question for the general conditions for a potential to exists for dissipative dynamical systems. They have shown that both ways of reasoning are equivalent, i.e., the potential obtained from Equ.(4) and the one from Equ.(7). They also study the concrete question of sufficient conditions for the existence of a potential, given a function $f(\vec{x})$. The method bases on solving the question of the integrability of a formal associated of a Hamilton systems to the Equ.(8). However, we demonstrate in the appendix, that their scheme is just the condition for certain diff-form to be closed [8], or equivalently, to posses an integrating factor. Also, we show below that for every dissipative systems there is a decomposition quantitatively similar to equation Eqn.(3). Remarkably, the question if a given dissipative systems is a gradient like systems also find relevance in mathematics. There are claims [10] that with high probability one will not be able to avert that question. However, the authors Chill and Fasarngova do not shrug off the importance of the question, on the contrary, they express hope by saying that they seem to see gradient systems every where and reaffirm the importance of the question by saying that they hail a gradient system whenever they see one. But their main inquires are how to show that a dissipative systems is a gradient systems and how to find the potential when one feels it exist?.
In this work we analyse this question[11], specifically for systems like:
$\beta_{c_{j}}\ddot{x}_{j}+\dot{x}_{j}+g_{j}(x_{1},\ldots,x_{N+1})+\zeta_{j}(t)=0;\hskip 14.22636ptj=1,\ldots,N+1$ (8)
These are in fact a sufficient general systems which contain interesting limits. In this system, the $x_{j}$ variables defines a vector in $\vec{x}\in R^{N+1}$, and the functions $\vec{g}=g_{j}(x_{1},\ldots,x_{N+1})$ defines the vector function $\vec{g}\in R^{N+1}$, these functions are assumed to be at least twice continuous differentiable, such that existence and uniqueness of classical solutions the ODE system is guarantee [12]. In addition, the last term represents a white noise term, ever present in physical systems, therefore we are dealing with stochastic differential equations. The parameters $\beta_{cj}$ is a measure of an inertia like concept in physics. In cases where the noise term can be neglected, justified for example at low enough temperatures, one can analyze the system by the flow properties of its associated first order system:
$\vec{\beta}_{c}\dot{\vec{x}}=\vec{\bar{x}}$
$\dot{\vec{\bar{x}}}=-\frac{1}{\beta_{c}}\vec{\bar{x}}-\vec{g}(\vec{x})$ (9)
where $\vec{\beta}_{c}=\beta_{1},....,\beta_{N+1}$, $\vec{x}=(x_{1},..,x_{N+1})$,  $\vec{\bar{x}}=(x_{N+1},..,x_{2(N+1)})$ are vectors $\in R^{N+1}$ [3][4]. If there exist a function , such that $\vec{\nabla}V(\vec{x})=\vec{g}(\vec{x})$, we say the system is gradient like, but note that It is not a gradient system. Here we are in the case of a the dynamical system which is phase space contracting like the systems treated in [6] and write for it for short: $\dot{x_{i}}=X_{i}$. For negligible inertia, one has the over-damped limit: $\dot{\vec{x}}=-\vec{g}(\vec{x})$. On the other hand, if the dissipative term (second term in Equ.(8)), or second term in the second equation in Equ.(9)) is negligible as compared with the inertial term, and there is a potential function $V(x_{1},..,x_{N+1})$, the systems is Hamiltonian with Hamilton function given by $\mathcal{H}_{P}(\vec{x},\vec{y})=\left(\beta_{c}\sum_{i=1}^{N+1}\frac{1}{2}y_{i}^{2}+V(x_{1},..,x_{N+1})\right)$:
$\beta_{c}\dot{\vec{x}}=\nabla_{\vec{x}}\mathcal{H}_{P}(\vec{x},\vec{y})$;   $\dot{\vec{y}}=-\nabla_{\vec{x}}\mathcal{H}_{P}(\vec{x},\vec{y})$.
II Differential Forms and the existence of a Potential.
Differential forms [13][14][15] have being used in expressing in a natural way fundamental theories like the Maxwell equations for electromagnetism or the Einstein general theory of relativity [16][17] (see [19]). Differential forms are basic for defining the symplectic structure of classical Hamiltonian mechanics[20]. For our purposes we introduce the essentials needed in the sequel, and adhere to the conviction in [18].
Let $x\in\mathcal{U}\subseteq R^{N+1}$, and $\omega(x)$ a function to the space of the p-linear alternating functions $\mathcal{A}_{p}(R^{N+1},R)$ , i.e., functions from $R^{p}\rightarrow R$, and the meaning of alternating just below. Naming the vector space of all differential forms by $\Omega^{n}_{p}(R^{N+1},R)$ which are of class $C^{n}$, $\omega(x)\in\Omega^{n}_{p}$ means $\omega(x,\xi_{1},....\xi{p})\in R$, and alternating in the variables $\xi_{j}$ means $\epsilon(\sigma)\omega(x,\xi_{\sigma(1)},....\xi_{\sigma(p)})=\omega(x,\xi_{1},....,\xi_{p})$, where $\epsilon(\sigma)$ denotes the signature of the permutation of the variables $\xi_{j}$[14]. Now naturally the question of differentiation appears, and the derived function[5] of $\omega(x,\xi_{1},....,\xi_{p})$ denoted by $\omega\prime$ is a functional from $U$ into the space $\mathcal{L}(R^{N+1};\mathcal{A}_{p}(R^{N+1},R)$, i.e., the space of functionals from $R^{N+1}$ into $\mathcal{A}_{p}(R^{N+1},R)$, i.e., $\omega\prime(x;\xi_{0})(\xi_{1},..,\xi_{p})\in R$, $(\xi_{0},\xi_{1},..,\xi_{p})\in R^{N+1}$. However, $\omega\prime(x;\xi_{0})(\xi_{1},..,\xi_{p})$ is not an alternating function of $(\xi_{0},\xi_{1},..,\xi_{p})$, it is an alternating function of the $(\xi_{j},j=1,..p$, and linear in the $(\xi_{0})$. To consistently stay within the multi-linear alternating class of functions, one apply the same procedure employed for the consistently definition of multiplication of two alternating functions, i.e., one ($f(x_{1},...,x_{p})\in R)$ in $\mathcal{A}_{p}(R^{N+1},R)$ and another ($g(x_{p+1},...,x_{p+q})\in R)$ in $\mathcal{A}_{p}(R^{N+1},R)$. The canonical form realized via a bilinear mapping $\phi_{p,q}:\mathcal{A}_{p,q}(R^{p},R)\rightarrow\mathcal{A}_{p+q}(R^{p},R)$, such that the product: $\it{h}:=f(x_{1},...,x_{p})g(x_{p+1},...,x_{p+q})\rightarrow\tilde{\it{h}}=\sum_{\sigma}\epsilon(\sigma)(\sigma\it{h})$, and the summations is over all permutations of the set $\{1,....p+q\}$ such that $\sigma(1)<,....,<\sigma(p)$, and $\sigma(p+1)<,....,<\sigma(p+q)$. Accordingly, summarizing we have then:
$x\in U,\hskip 8.5359pt\omega(x)\Rightarrow\mathcal{A}_{p}(R^{N+1},R)$
$\omega^{\prime}(x):\mathcal{U}\Rightarrow\mathcal{L}(R^{N+1};\mathcal{A}_{p}(R^{N+1},R))$
$\omega^{\prime}(x;\xi_{0})(\xi_{1},..,\xi_{p})\in\mathcal{A}_{1,p}(R^{N+1},R)\stackrel{{\scriptstyle d\omega}}{{\longrightarrow}}\mathcal{A}_{1+p}(R^{N+1},R)$
$d\omega(x;\xi_{0},\xi_{1},..,\xi_{p})=\sum_{i=1}^{p}(-1)^{i}\omega^{\prime}(x;\xi_{i})(\xi_{1},..,\tilde{\xi}_{i},.....\xi_{p})$
$d(d\omega(x;\xi_{0},\xi_{1},..,\xi_{p}))=0$ (10)
The tilde over $\xi_{i}$ means it is omitted from the variables list. This explicit expression for the d-operator $d\omega$ is a result of the definition above for product of differential forms. Now we must consider another operation on differential forms, i.e., the change of variable procedure[18]. Given a mapping $\phi\in C^{n+1}$ from a subspace $U^{\prime}\in R^{N+1}$ to another $U\in R^{N+1}$, and given that one has differential p-forms defined in $U$, the mapping permits to define differential forms on $U^{\prime}$. We assume that the inverse mapping, $\phi^{-1}$ exist and $\in C^{n+1}$:
$\vec{x}\in U,\hskip 8.5359pt\omega(x)\Rightarrow\mathcal{A}_{p}(R^{N+1},R)$
$\phi:U^{\prime}\rightarrow U,\hskip 5.69046pt\vec{y}\in U^{\prime},\hskip 2.84544ptx_{i}=\phi_{i}(y_{1},..,y_{N+1}),\hskip 8.5359pt\phi^{\star}(\omega):U^{\prime}\stackrel{{\scriptstyle\phi^{\star}}}{{\longrightarrow}}\mathcal{A}_{p}(R^{N+1},R)$
$(\eta_{1},..,\eta_{p})\in R^{N+1},\hskip 8.5359pt(\phi^{\star}(\omega))(y;\eta_{1},..,\eta_{p}):=\omega(\phi(y);\phi^{\prime}(y)\cdot\eta_{1}),.....,\phi^{\prime}(y)\cdot\eta_{p})$
$d(\phi^{\star}(\omega))=\phi^{\star}(d\omega)$ (11)
The change of coordinate is a linear mapping in $\Omega^{n}_{p}(R^{N+1},R)$, the last line express the core of the consistency of the theory, the value of the $d\omega$ is independent of the coordinates systems, i.e., invariant under change of variables[15]. Another remarkably aspect of the theory is Poincare $\grave{}s$ theorem [21]. Assume from now on that the set $U$ is starlike with respect one of its points say the point a, if the for all x, $\in U$, the segment (1t)a+tx(1-t)a+tx, $0\leq t\leq 1$ is contained in U [16], in particular a convex set is starlike with respect to any of its points. One can take in full generality $a=0$, the zero of the vector space $R^{N+1}$. We now consider particular mapping that was first proposed in [23] and define the operator $k(\omega)$, that acts on differential forms:
$(\xi_{1},..,\xi_{p-1})\rightarrow\omega(tx;x,\xi_{1},..,\xi_{p-1}),\hskip 5.69046pttx\in U$
$k(\omega)(x,\xi_{1},..,\xi_{p-1})=\int\limits_{0}^{1}t^{p-1}\omega(tx;x,\xi_{1},..,\xi_{p-1})dt$
$k:\Omega^{n}_{p}(R^{N+1},R)\rightarrow\Omega^{n}_{p-1}(R^{N+1},R)$
The integral is understood in the usual Riemann sense. The mapping is particular in which the argument of the differential form $\omega$ for which it is defined in $U$ is the same as one of the first vector where it must be evaluated to obtain finally its value $\in R$. Therefore, the operator $k(\omega)$ is alternating in the variables $(\xi_{1},..,\xi_{p-1})$. In addition, the differentiability assumptions on $(\omega)$ permits to differentiate under the integral sign, then the important result is obtained [14], [16]:
$d(k(\omega))+k(d\omega)=\omega$ (13)
[…]
$\omega:=\mathcal{G}=\sum_{j=1}^{N+1}g_{j}(x_{1},\ldots,x_{N+1})dx_{j};\hskip 14.22636pt\mathcal{G}\in\Omega^{n}_{1}(R^{N+1},R),$
$\mathcal{G}(\vec{x},\vec{\Xi})=\sum_{j=1}^{N+1}g_{j}(x_{1},\ldots,x_{N+1})\xi_{j}$
$d\mathcal{G}(\vec{x},\xi_{1},\xi_{2})=(\omega^{\prime}(\vec{x};\xi_{1}))(\xi_{2})-(\omega^{\prime}(\vec{x};\xi_{2}))(\xi_{1})=0\hskip 14.22636pt\Rightarrow$
$\left[g_{j}(\vec{x}),\partial{x_{l}}\right]:=\frac{\partial{g_{j}(\vec{x})}}{\partial{x_{l}}}=[g_{j}(\vec{x}),\partial{x_{l}}]^{T}=[g_{l}(\vec{x}),\partial{x_{j}}]$ (14)
The last conditions says that the matrix $g_{j}(\vec{x}),\partial{x_{l}}]_{j,l}$ of the partial derivatives is symmetric. In the sequel we prove previous statement for our $p=1$ diff-form. From Equ.(14) and the definition Euq.(12) we get:
$k(d\omega)((\vec{x})\cdot\xi_{1}=\int\limits_{0}^{1}t[(\omega^{\prime}(\vec{tx};\vec{x})\cdot\xi_{1})-(\omega^{\prime}(\vec{tx};\xi_{1})\cdot\vec{x})]dt=$
$\int\limits_{0}^{1}t\sum_{i,j;i\neq j}[\frac{\partial{g_{i}(\vec{t\vec{x}})}}{\partial{x_{j}}}-\frac{\partial{g_{j}(\vec{t\vec{x}})}}{\partial{x_{i}}}](x_{i}\xi^{j}_{1}-\xi^{j}_{1}x_{i})dt$
$d(k(\omega))(\vec{x})\cdot\xi_{1}=\int\limits_{0}^{1}[t(\omega^{\prime}(\vec{tx};\xi_{1})\cdot\vec{x})+\omega(\vec{tx})\cdot\xi_{1}]dt\hskip 39.83368pt$
Then, we have the following facts:
$\frac{\partial{g_{i}(\vec{t\vec{x}})}}{\partial{x_{j}}}=\frac{\partial{g_{j}(\vec{t\vec{x}})}}{\partial{x_{i}}}\iff d\omega=0\iff[(\omega^{\prime}(\vec{tx};\vec{x})\cdot\xi_{1})=(\omega^{\prime}(\vec{tx};\xi_{1})\cdot\vec{x})]\Rightarrow$
$t(\omega^{\prime}(\vec{tx};\xi_{1})\cdot\vec{x})+\omega(\vec{tx};\vec{x})\cdot\xi_{1}=(t\omega^{\prime}(\vec{tx};\vec{x})+\omega(\vec{tx};\vec{x}))\cdot\xi_{1}=\frac{d(t\omega(t\vec{x}))}{dt}\cdot\xi_{1}\Rightarrow$
$d(k(\omega))(\vec{x})=\int\limits_{0}^{1}\frac{d(t\omega(t\vec{x}))}{dt}=\omega(\vec{x})$
See [27]. Although we are dealing with differential forms, the last relation can be considered as an equality of functions in $\mathcal{A}_{1}(R^{N+1},R)$. This procedure is central to our reasoning below where we are pursuing the functions defining the p=1 diff=form. As a result, we have that:
$k(\omega)(\vec{x})=\int\limits_{0}^{1}\omega(\vec{tx};\vec{x})dt=\int\limits_{x_{0}=0}^{\vec{x}}\omega(\vec{x})dx=\sum^{N}_{i=1}\int\limits_{0}^{1}x_{i}g_{i}(t\vec{x})dt$
$\Rightarrow\frac{\partial{k(\omega)(\vec{x})}}{\partial{x_{i}}}=g_{i}(\vec{x})$
the integral is along a linear segment from $x_{0}$ to $x$ [14]. We name the second line in previous equation the consistency condition. Now, $k(\omega)(\vec{x})$ defines a function in $U$, and the integral of $\omega(\vec{x})$ around any class $C^{1}$ loop $\gamma$ is zero[14], in fact, $d\mathcal{G}=0\iff\int\limits_{\gamma}\omega(\vec{x})=0$. Equ.(17) defines the potential. When $d\mathcal{G}\neq 0$, i.e., $\mathcal{G}$ is not closed, the first term in Eqn.(13) defines a closed form, ($dd(k(\omega))=0)$, and and a non-coed form called the residual part $k(d\mathcal{G})$. In other words, the lack of the nice property of the equality of the cross derivatives that straightforwardly conducted us to the potential function, now makes the a residual part according to Eqn.(15). In either case, the matrix of the partial derivatives $\left[g_{j}(\vec{x}),\partial{x_{l}}\right]$ is the main object to analyze. No coordinate transformation bring us outside this relation, whose demonstration is now straightforward as follows: Consider the sum of both terms in Eqn.(15),
$k(d\omega)((\vec{x})\cdot\xi_{1}+d(k(\omega))(\vec{x})\cdot\xi_{1}=\int\limits_{0}^{1}t[(\omega^{\prime}(\vec{tx};\vec{x})\cdot\xi_{1})-(\omega^{\prime}(\vec{tx};\xi_{1})\cdot\vec{x})]dt+$
$\int\limits_{0}^{1}[t(\omega^{\prime}(\vec{tx};\xi_{1})\cdot\vec{x})+\omega(\vec{tx})\cdot\xi_{1}]dt=\int\limits_{0}^{1}[t(\omega^{\prime}(\vec{tx})\cdot\vec{x})+\omega(\vec{tx})dt]\cdot\xi_{1}=$
$\int\limits_{0}^{1}\frac{d(t\omega(t\vec{x}))}{dt}\cdot\xi_{1}=\omega(\vec{x})\cdot\xi_{1}=\mathcal{G}(\vec{x},\vec{\Xi});\hskip 56.9055pt$
i.e., the cancelation of the first term in the third line and second term in first line in Equ.(15) permits the same manipulations given in the second line in Equ.(16) [22][7]. From Equ.(11), the non-closed condition can not be cure by any change of variables. Then, from Eqn.(16), given a dissipative systems the first thing to check pursuing the existence of a potential function is the equality of the cross-derivatives of the associated p=1 diff-form. When this is not the case, the following ideas are pertinent.
II.1 Interpretation of the dynamics in term of Differential Forms
Now, a central point in our approach is to interpret the differential equation Equ.(9) in terms of differential forms, fort his purpose we define:
$\dot{\vec{y}}=Y(\vec{y})\Rightarrow\Psi:=d\vec{y}-Y(\vec{y})dt$ (19)
The p=1 differential form $\Psi$, defined in a set $U\subseteq R^{n}XR$. When $Y_{i}(\vec{y})$ $\in C^{1}(\vec{y})$, Frobenious theorem [14] establish the existence of a unique solution $\vec{y}_{s}(t,\vec{y})$, the one that annihilates the differential form $\Psi$, i.e., $\dot{\vec{y}}_{s}(t,\vec{y})dt-Y(\vec{y}_{s})dt=0$, and so therefore, by considering equality between elements in $\Omega^{n}_{1}(R^{N+1},R)$, we have within the differential forms formalism the validity of following relation[25](see Eqn.(9)):
$\beta_{c}d\vec{x}=\vec{y}dt\hskip 332.89716pt$
$d\vec{y}=-\frac{1}{\beta_{c}}\vec{y}dt-[d(k(\mathcal{G}))(\vec{y})+k(d\mathcal{G})(\vec{y})]dt;\hskip 163.03416pt$
$d(k(\mathcal{G}))(\vec{y})+k(d\mathcal{G})(\vec{y})=\int\limits_{0}^{1}t(\mathcal{G}^{\prime}(\vec{tx})\cdot\vec{x})dt+\int\limits_{0}^{1}\mathcal{G}(\vec{tx})dt\hskip 113.81102pt$
The whole bracket make sense as a function, i.e., the considered equality of elements in $\Omega^{n}_{1}(R^{N+1},R)$ works, i.e., we always have in mind the canonical representation of diff-form, and for conclusions only the defining function of the diff-form matters. We conclude, by comparing Equ.(20) with Equ.(3), this the most general possible case for any systems defined by Equ.(9), and that the orthogonality condition is a particular one. Any suitable change of variables $\phi_{i}(y_{1},..,y_{N+1})$ that express the same SOD in other coordinates [30] leaves invariant the structure of Equ.(20). As an example, we take the Lorenz equations[3]. With $\vec{\mathbf{y}}=(x,y,z)$$\in R^{3}$, the vector function $\vec{\mathbf{G}}=(g_{1}(\vec{\mathbf{y}}),g_{2}(\vec{\mathbf{y}}),g_{3}(\vec{\mathbf{y}}))$, with ($g_{1}(\vec{\mathbf{y}})=\sigma(y-x),\hskip 5.69046ptg_{2}(\vec{\mathbf{y}})=\rho x-y-xz,\hskip 5.69046ptg_{3}(\vec{\mathbf{y}})=-\beta z+xy)$, and the parameters $\sigma,\rho,\beta$ are >0>0. For the dissipative system $\dot{\vec{\mathbf{y}}}=\vec{\mathbf{G}}$, one easily verifies that the matrix function of the partial derivatives is not symmetric, in fact one finds $\partial g_{i}(\vec{\mathbf{y}})/\partial\mathbf{y}_{j}\neq\partial g_{j}(\vec{\mathbf{y}})/\partial\mathbf{y}_{i}$, $1\leq i,j\leq 3$. Therefore, for the associated $p=1$ differential form $\mathcal{G}(\vec{\mathbf{\mathbf{y}}})$(Equ.(14)), $d\mathcal{G}(\vec{\mathbf{\mathbf{y}}})\not=0$, and therefore, there is not a potential in the sense of Equ.(2,17). In addition, we have $\vec{\mathbf{G}}\cdot\nabla\times\vec{\mathbf{G}}\neq 0$ and there is no sense in trying the Graham scheme (see appendix, Eqn.(35)), [8][9], notwithstanding, the universal relation applies(see[28], theorem 5-3.1, where the operator k is denoted by $\mathcal{H}$, the Homotopy operator, see also theorem 5-6.1, and the corollaries 5-6.1-3):
$\omega=\omega_{exact}+\omega_{antiexact}:=d(k(\mathcal{G}))(\vec{\mathbf{y}})+k(d\mathcal{G})(\vec{\mathbf{y}})\Rightarrow$
$d(\omega)=d(d(k(\mathcal{G}))(\vec{\mathbf{y}}))+d(k(d\mathcal{G})(\vec{\mathbf{y}}))=d(k(d\mathcal{G})(\vec{\mathbf{y}}))=d\mathcal{G}(\vec{\mathbf{y}})$
$k(\omega)=k(d(k(\mathcal{G}))(\vec{\mathbf{y}}))+k(k(d\mathcal{G})(\vec{\mathbf{y}}))=k(d(k(\mathcal{G}))(\vec{\mathbf{y}}))=k(\mathcal{G})$
The descomposition of $\omega$ is such, that the first term is in the kernel of the operator d, and the second term in the kernel of the operator k. From visual comparison of previous equation with Equ.(3), we define in full generality, the first term $\omega_{exact}:=dk(\mathcal{G}))(\vec{y})$ as the potential part, it is a closed form ($dd(k(\mathcal{G}))(\vec{\mathbf{y}})=0)$ and therefore ($d(k(\mathcal{G}))(\vec{\mathbf{y}})$ is an exact form, therefore there exist a function $\mathcal{G}_{pot}:=k(\mathcal{G}))(\vec{\mathbf{y}})$, such that $d\mathcal{G}_{pot}=\omega_{exact}$. In analogy, we name the second term, $\omega_{antiexact}:=k(d\mathcal{G})(\vec{y})$ as the residual part, it is invariant under the action of the operator $(dk)$ and belongs to the kernel of the operator k [26], $k(k(d\mathcal{G}))=0$ , and therefore there is a p=1+1 diff-form, namely and explicit $d\mathcal{G}$. In fact, [28], $\omega_{exact}$ and $\omega_{antiexact}$ belong to sets whose intersection is the null element in the set of the $p=1$ differential forms, this property can be considered as an orthogonality like relation, and this means, the representation through exact and antiexact forms is unique. For the Lorentz system, es is an easy task to calculate $\mathcal{G}_{exact}$ and $\mathcal{G}_{antiexact}$. But the message is general, to any dissipative system one can associate a differential form $\mathcal{G}$, defined in Eqn.(14), from which a potential function and a non-potential function also called the residual part, can be derived ($\omega_{exact}$ and $\omega_{antiexact}$). This generalizes the approach defined by Eqn.(3). It it happens that the associated form is closed, the system is a gradient or a gradient like system. However, there are claims [10], that every dissipative systems should be a gradient system. In the next section we duel on this claim.
III A Generalized change of Variables
As explained above, the central objet of the discussions is the matrix of the partial derivatives $\left[g_{j}(\vec{x}),\partial{x_{l}}\right]$. If it is a symmetry matrix, the associated differential form to the dissipative systems $\mathcal{G}$ is closed, and therefore there is a potential. It is not a symmetry matrix, there is a potential part and an additional form term which is not exact. This state of affairs is invariant within the change of variables procedure , Eqn.(11). One has to resort to a non-linear kind of transformation as possible road map of remedy. that is, we want to create from one system of ODE, another one through this non-linear transformation by maintaining the equivalence of the two. In this new ODE, the closed condition on an associated p=1 differential form is imposed, and the solution derived from it, is used to find the potential. We introduce a generalized change of variables $\phi^{\star\star}$(compare Eqn.(11)): given a a function $\phi_{i}(y_{1},..,y_{N+1})$ $\in C^{n}(\vec{y})$, for our case of interest, $p=1$, and such that the inverse function $\phi^{-1}(\vec{x})$ exist and ist differentiable:
$\vec{x}\in U,\omega\in\Omega^{n+1}_{p=1}(U,R),\hskip 8.5359pt\omega(x)=\sum_{j=1}^{N+1}g_{j}(\vec{x})dx_{j}\in\mathcal{A}_{p=1}(R^{N+1},R),\hskip 28.45274pt$
$\phi:U^{\prime}\rightarrow U,\hskip 5.69046pt\vec{y}\in U^{\prime},\hskip 11.38092ptx_{i}=\phi_{i}(y_{1},..,y_{N+1}));\hskip 128.0374pt$
$\phi^{\ast}(\omega)\in\Omega^{n+1}_{p=1}(U^{\prime},R),\hskip 5.69046pt\phi^{\ast}(\omega)=\sum_{j=1}^{N+1}g_{j}(\phi(\vec{y}))d\phi_{j}(y_{1},..,y_{N+1})=\sum_{j,i=1}^{N+1}g_{j}(\phi(\vec{y}))\frac{d\phi_{j}(y_{1},..,y_{N+1})}{dy_{i}}{dy_{i}}$
$\phi^{\star\star}(\omega)(\vec{y}):=\phi(\vec{g}))(\phi^{-1}(\vec{y}))=\sum_{j}^{N+1}\phi_{j}(g_{j})(\phi^{-1}(\vec{y})){dy_{j}}\hskip 56.9055pt$
i.e., $\phi^{\ast\ast}(\omega)$ is a $p=1$ form from $U^{\prime}\rightarrow\mathcal{A}_{p=1}(R^{N+1},R)$ in which the functions defining the canonical form are transformed by the function $\phi$ which also defines the coordinate transformation, in contrast to $\phi^{\ast}(\omega)$, $\phi^{\ast\ast}(\omega)$ is nonlinear in $\Omega^{N+1}_{p=1}(R^{N+1},R)$. With the previous definition we transform Equ.(8) to the new variables $\vec{y}$, without noise part [29]:
$\dot{x_{i}}=X_{i}(\vec{x}))\Rightarrow\dot{y_{i}}=Y_{i}(\vec{y})\hskip 85.35826pt$
$\beta_{c}\ddot{y_{i}}+\ddot{y_{i}}+f_{i}(\phi^{-1}(y_{1}),\ldots,\phi^{-1}(y_{N+1}))=0;\hskip 14.22636ptj=1,\ldots,N+1$
$f_{i}(\phi^{-1}(y_{1}),\ldots,\phi^{-1}(y_{N+1})):=\phi_{i}(g_{1}(\phi^{-1}\vec{y}),\ldots,g_{N+1}(\phi^{-1}\vec{y}))$
$\mathcal{F}:=\sum_{i=1}^{N+1}f_{i}(\phi^{-1}(y_{1}),\ldots,\phi^{-1}(y_{N+1}))dy_{i}\hskip 71.13188pt$
$\phi_{i}(\vec{y})=\sum_{j}\mathcal{D}_{i,j}(\vec{\mathbf{y}})y_{j}\hskip 113.81102pt$ (23)
[…]
$\vec{x}=\mathcal{D}(\vec{y})\vec{y}\hskip 5.69046pt\Rightarrow\vec{y(\vec{x})}=\mathcal{D}^{-1}(\vec{y}(\vec{x})\vec{x}\hskip 2.84544pt\Rightarrow\frac{\partial g_{j}(\vec{y}_{j})}{\partial x_{k}}=\sum_{p}\frac{\partial g_{j}(\vec{y}_{j})}{\partial y_{p}}\frac{\partial(\vec{y}_{p})}{\partial x_{k}}=$
$\sum_{p}\frac{\partial g_{j}(\vec{y}_{j})}{\partial y_{p}}(\mathcal{D}^{-1}_{p,k}(\vec{{y}}(\vec{x}))+\sum_{l}\frac{\partial\mathcal{D}_{p,l}(\vec{{y}}(\vec{x}))}{\partial x_{p}}\cdot x_{l})\hskip 1.42271pt\Rightarrow$
$\left[\frac{\partial f_{i}(\vec{y})}{\partial x_{k}}\right]=\left[\frac{\partial f_{i}(\vec{y})}{\partial x_{k}}\right]^{T}=\left[\frac{\partial f_{k}(\vec{y})}{\partial x_{i}}\right]=\left(\sum_{j}\mathcal{D}_{i,j}(\vec{y}(\vec{x}))\hskip 1.42271pt\sum_{p}\frac{\partial g_{j}(\vec{y}(\vec{x}))}{\partial y_{p}}\hskip 1.42271pt\mathcal{D}^{-1}_{p,k}(\vec{y}(\vec{x}))\right)+$
$\left(\sum_{j}\mathcal{D}_{i,j}(\vec{y}(\vec{x}))\sum_{p}\frac{\partial g_{j}(\vec{y}(\vec{x}))}{\partial y_{p}}\sum_{l}(\frac{\partial\mathcal{D}_{p,l}}{\partial x_{k}})\cdot x_{l}\right)+\left(\sum_{j}\frac{\partial\mathcal{D}_{i,j}(\vec{y}(\vec{x}))}{\partial x_{k}}g_{j}(\vec{y}(\vec{x}))\right)$
Or equivalently with the definitions: $g_{j}:=g_{j}(\vec{y}(\vec{x}))$, $g_{j,\partial y_{k}}:=\frac{\partial g_{j}(\vec{y}(\vec{x}))}{\partial y_{k}}$, , and   $\mathcal{D}_{i,j,\partial x_{k}}:=\frac{\partial\mathcal{D}_{i,j}(\vec{y}(\vec{x}))}{\partial x_{k}}$, and implicit sum over repeated indices :
$\left(\left[\mathcal{D}_{i,j}g_{j,\partial y_{p}}\mathcal{D}^{-1}_{p,k}\right]-\left[\mathcal{D}_{i,j}g_{j,\partial y_{k}}\mathcal{D}^{-1}_{p,k}\right]^{T}\right)+\hskip 113.81102pt$
$\left(\left[\mathcal{D}_{i,j}g_{j,\partial y_{p}}\mathcal{D}_{p,l,\partial x_{K}}\cdot x_{l}\right]-\left[\mathcal{D}_{i,j}g_{j,\partial y_{p}}\mathcal{D}_{p,l,\partial x_{K}}\cdot x_{l}\right]^{T}\right)+\hskip 56.9055pt$
$\left(\left[\mathcal{D}_{i,j,\partial x_{k}}\cdot g_{j}\right]-\left[\mathcal{D}_{i,j,\partial x_{k}}\cdot g_{j}\right]^{T}\right)=0\hskip 56.9055pt$
The consistency condition is given by:
$k(\mathcal{F})(\vec{x})=\int\limits_{0}^{1}\mathcal{F}(\vec{tx};\vec{x})dt=\int\limits_{x_{0}=0}^{\vec{x}}\mathcal{F}(\vec{x})dx=\sum^{N}_{i=1}\int\limits_{0}^{1}x_{i}f_{i}(t\vec{x})dt:\hskip 5.69046pt\frac{\partial{k(\mathcal{F}))(\vec{x})}}{\partial{x_{i}}}=f_{i}(\vec{x})$
The nonlinear differential equation for $\mathcal{D}_{i,j}(\vec{y}(\vec{x}))$ subject to the consistency condition may be remarkably awkward for some systems as we have already checked , we treat therefore first the simplest cases where $\mathcal{D}_{i,j}(\vec{\mathbf{y}})$ is a constant matrix $\mathcal{D}_{i,j}$, i.e., applicable when the matrix of the partial derivatives $g_{j,\partial y_{k}}$ is a constant matrix. In these cases only the first term in parenthesis in Eqn.(26) is non zero. One obtains the following matrix equation:
$x_{i}:=\sum_{j=1}^{N+1}\mathcal{D}_{i,j}y_{j},\hskip 5.69046ptf_{i}(\phi^{-1}(\vec{x})):=\sum_{j=1}^{N+1}\mathcal{D}_{i,j}g_{j}(\phi^{-1}(\vec{x})))$
[…]
If the form $\mathcal{G}$ where closed, then the solution is the identity matrix $\mathcal{D}=I$. The last equation constitutes a necessary condition to be fulfilled by the solution of $g,\partial{y}]$ as a function of $\mathcal{D}$, which is obtained from the Eqn.(17). Therefore, with the consistency condition Eqn.(27) one obtains:
$k(\omega)(\vec{x})=\int\limits_{0}^{1}\omega(\vec{tx};\vec{x})dt=\int\limits_{x_{0}=0}^{\vec{x(y)}}\omega(\vec{x})dx\Rightarrow\mathcal{D}(\mathcal{D}^{T})^{-1}=[g,\partial{y}]^{T}$ (29)
The solution fulfills the necessary condition Eqn.(28). As the matrix $g,\partial{y}]$ is constant, a straightforward way to get the last relation is to use the ansatz for the potential $U=\sum_{i}\frac{1}{2}\phi_{i}(\vec{x})^{2}$ with $\phi_{i}(\vec{x})=f_{i}(\vec{x})$ and to apply on $U$ the d-operator and then dd.
IV Application of the generalized change of variables concept.
IV.1 Two-dimensional Josephson Junctions Arrays:
The resistive behavior of a two dimensional Josephson array with a given frustration parameter (the ratio of the perpendicular magnetic field to the flux quantum per plaquette ($f=\frac{M}{N}[\Phi_{0}]$, where $M,N$ are natural numbers , with $\Phi_{0}=hc/2e$ ), have been a matter of intense research [35][37][38], see page 21 in [39]. In this case, primary there are phase variables for which there are functions $\Phi_{k}(y)$ which express their dependence on the variables $y$, and an auxiliary matrix $\omega$. Then we have for any $(M,N)$ for the functions appearing in Eqn.(8) are given by Eq.(3) in [35], where $\forall j,\beta_{cj}=\beta_{c}$. Fist, accordingly, one associates a differential form to the dissipative system and checks the cross derivatives to find that $\frac{\partial g_{j}(\vec{y}(\vec{x}))}{\partial y_{k}}\neq\frac{\partial g_{k}(\vec{y}(\vec{x}))}{\partial y_{j}}$, i.e., the associated diff-form is not closed. Then the second step is carried out and we obtain with $\vec{x}=\mathcal{D}\vec{y}$, and $a_{ik}:=\frac{1}{2}\sum_{j=1}^{N+1}D_{ij}\omega_{jk}$, the expression for the $f_{i}(\vec{x})$(Eqt.(5) in [35]), and so finally the specific equation for the matrix $\mathcal{D}$ is derived from Equ.(28):
$\mathcal{D}^{T}\mathcal{D}[\omega\Phi_{k,\partial{y}}]=[\omega\Phi_{k,\partial{y}}]^{T}\mathcal{D}^{T}\mathcal{D}\hskip 5.69046pt\Rightarrow\frac{1}{2}\mathcal{D}^{T}\mathcal{D}=\Phi_{\partial{y},k}\omega.$ (30)
With the help of Equ.(30), the potential can be calculated in principle for any $M,N$ (see Eqn.(6) and Eqn.(7) in [35]), and in particular, we have calculated the potential for two important cases, namely, $f=\frac{1}{2},\frac{1}{3}$ [35],[36]. The authors in [40], [41],[42] among many others , hail the importance of these two cases. There are noise terms associated with these systems, and is relevant to say that theses terms transforms differently in the new systems of coordinates defined by the matrix $\mathcal{D}$, such that, the intensity of the noise is bigger for some variables (for details for $f=\frac{1}{2}$ see [35]). This a matter with physical consequences when one consider the whole stochastic behavior of the system (this issue is under study). Furthermore, they find the case $f=\frac{2}{5}$ is for some unknowns reasons a case that emulate in a universal form cases near incommensurability where $f=(3-\sqrt{5})/2$. The matrix $\mathcal{D}$ for $f=\frac{2}{5}$ is a complicated object versions of which we have already derived. However, a static disorder due to a non-vanishing variance in the Josephson parameters makes the SDE also a quenched disordered system[44]. Summarizing, the dynamic equations for two-dimensional Josephson arrays under a perpendicular magnetic field are gradient like systems with the particularity that the potential are not bounded from below being tilted washboard functions.
IV.2 Examples with Josephson Junctions Circuits.
There are many models of Josephson circuits [33]. We use Equ.(28) to obtain the potential for the Josephson junction circuit defined in [34]. We write their equations (2.42.7)(2.4-2.7) in the form of Equ.(8) where thermal noise is neglected for the variables $\vec{y}=(y_{1}=y,y_{2}=\delta,y_{3}=\zeta)$, and the three functions $\vec{g}=g_{1}(\vec{y}),g_{2}(\vec{y}),g_{3}(\vec{y})$.
$\beta_{c}\ddot{\delta}=-r\dot{\delta}+(\imath-\sin{\delta}-\zeta)$
$\frac{1}{\beta_{L}}\dot{\zeta}=-\zeta+y\hskip 69.70915pt$
The last equations give rise to a dissipative system in the form of Equ.(9):
$\hskip 85.35826pt\dot{\delta}=y=:g_{1}(y)\hskip 150.79968pt$
$\dot{y}=-\frac{r}{\beta_{c}}y+\frac{1}{\beta_{c}}(\imath-\sin{\delta}-\zeta)=:g_{2}(y,\delta,\zeta)\hskip 31.2982pt$
$\dot{\zeta}=-\frac{1}{\beta_{L}}\zeta+\frac{1}{\beta_{L}}y=:g_{3}(y,\zeta)\hskip 92.47145pt$
If we can find a potential it constitutes a gradient system. One first verifies that $\partial g_{i}(\vec{\mathbf{y}})/\partial\mathbf{y}_{j}\neq\partial g_{j}(\vec{\mathbf{y}})/\partial\mathbf{y}_{i}$, $1\leq i,j\leq 3$, and therefore, for the associated $p=1$ differential form $\mathcal{G}(\vec{\mathbf{\mathbf{y}}})$(Equ.(14)) $d\mathcal{G}(\vec{\mathbf{\mathbf{y}}})\not=0$, and therefore, there is not a potential in the sense of Equ.(2,17). In this form, the Josephson junction circuit was numerically simulated for particular values of the parameters $(\imath,r,\beta_{c},\beta_{L})$[34]. We find the potential in two steps: First we study the linear case for small value of $\delta<<1$, and obtain accordingly previous explanations the following matrix equation:
$\mathcal{D}(\mathcal{D}^{-1})^{T}=[g_{i}(\vec{y}),\partial{y_{j}}]^{T}$ (33)
We then obtain the following matrix equation for the case under study, i.e., for the $3X3$ matrix $\mathcal{D}$:
$g_{i}(\vec{y}),\partial{y_{j}}]=\left(\begin{array}[]{llcl}1&0&0\\ \frac{-r}{\beta_{c}}&\frac{-1}{\beta_{c}}&\frac{-1}{\beta{c}}\\ \frac{1}{\beta{L}}&0&\frac{-1}{\beta_{L}}\\ \end{array}\right)=\mathcal{D}^{T}(\mathcal{D}^{-1})$
This matrix equations defines $9$ polynomial relations in each case for the $9$ constituents components $d_{ij}$ of $\mathcal{D}$, $1\leq i\leq 3,1\leq j\leq 1$. No analytical solution seems achievable, and only a systematic numerical study is possible.
From Eqn.(14) the following relations holds for $(\phi(\vec{y})=\mathcal{D})$:
$\forall i\neq j,\hskip 5.69046pt\frac{\partial f_{i}(\mathcal{D}^{-1}(\vec{x}))}{\partial x_{j}}=\frac{\partial f_{j}(\mathcal{D}^{-1}(\vec{x}))}{\partial x_{i}}$ (34)
and therefore, for p=1-diff-form $\mathcal{F}:=\sum_{i=1}^{N+1}f_{i}(\mathcal{D}^{-1}(x_{1}),\ldots,\mathcal{D}^{-1}(x_{N+1}))dx_{i}$ is closed and exact. The potential is $k(\mathcal{F}(\vec{x}))=\int\limits_{x_{0}=0}^{\vec{x}}\mathcal{F}(\vec{x})dx$. After that, we put a $cos(\delta)$ in place of the variable $\delta$ alone. Notwithstanding, the road map is to apply the procedure given in Eqn.(26) and Eqn.(27). In a similar way, the model of the circuit [43] can be analyzed, and is actually a matter under study.
V Summary and Perspectives.
In this work, two relevant results are obtained. On the one hand, from the theory of differential forms, we show that a p=1 differential form can be associated with a dissipative systems. This diff-form may be closed or not. If it is a closed form then we can associate to the dissipative system a potential. it constitutes a gradient system or a gradient like system. If it is no close, (Eqn.(14), there exist no coordinate transformation that could change this property. The property of been non-closed is invariant under coordinate transformation. Through the introduction of the Homotopic or k-operator, in combination with the d-operator, in either case, one can decompose the diff-form in an exact pat and an antiexact part. This decomposition is universal valid, and express that to any dissipative systems, its dynamics is given by a potential like part and a non-potential term. On the other hand, If it is not a closed form, we proposed in this work , a nonlinear matching procedure by ”rotating” both, the vector function $\mathcal{G}(\vec{y})$ and the coordinate system $\vec{y}$, simultaneously under the transformation $\phi(\vec{y})$ and imposing for the new p=1 diff-form the necessary condition for being closed. This is the generalized change of variables (Eqn.(22) which we implement through the introduction of a matrix function (Eqn.(23). For the p=1 diff-form constructed (Eqn.(23), we derive an equation (Eqn.(26)) whose results give the equation defining this transformation, together with the consistency condition (Eqn.(27). It constitutes our main result. We claim that the same conditions that guaranties the existence of a solution of the dynamical systems (Eqn.(9)), also guaranties the solution a Eqn.(26). For some cases, the transformation can be defined through a constant matrix, and an analytical solution can be found, i.e., the case of two dimensional Josephson arrays under a magnetic field [35], [36]. For the cases in [43], and [34], it appears to us that only a numerical study is feasible. These results bring us to the main purpose of this work, namely, to answer the question, is a dissipative systems always a gradient system ?. The examples discussed above, show certainly, that whereas the differential form formalism say there is not a potential, our approach finds that indeed it exist, in the context of the generalized change of variable concept. For the time being, we can only say for sure, that for the kind of systems where Eqt.(30) apply, one can find a potential. Also, for systems where Eqt.(33) applies, although the resulting algebraic equations appears to be accessible only through pure numerical simulations. And in general, depending on each particular system, one has to figure out a strategy for find the generalized transformation defined in Then we have the natural question: given a particular dissipative systems there exist such a transformation $\phi(\vec{y})$, as defined in Eqn.(23), , and such that it satisfies Eqn.(25), Eqn.(26), and Eqn.(27). This procedure may be a difficult one, but even in principle, all dissipative systems are gradient systems , or gradient like systems. For extended systems represented by an evolution equation like a parabolic partial differential equations, the corresponding Banach space is of infinite dimension, and the associated operators are better treated in a Hilbert space with a combination of the concepts of differential forms and functional analysis. This matter is under investigation. Summarizing, our approach to the problem to the existing of a potential for a dissipative dynamical systems provides a clear mathematical set up for finding the potential, and confirms the equality of the class of dissipative systems and the class of gradient systems.
VI Apendix
VI.1 Discussion of Graham Theory in the light of Differential forms.
In Graham scheme the dissipative model is given from the upset as in Eqn.(4). Their claim is referred to the existence of a potential with respect to a a certain given positive semi-definite matrix $\mathcal{S}$. In the context of this particular class of continuous dissipative systems , they showed that the existence of the potential is not a generic property. It consist in a decomposition (assuming possible) of a given function that govern the dynamics, i.e., $\dot{\vec{\tilde{x}}}=f(\vec{x})$, in the form: $f(\vec{x})=\mathbf{D}\nabla\mathbf{V}=(\mathbf{D}^{symm}+\mathbf{D}^{skew})\nabla\mathbf{V}$ in a part proportional to a positive definite symmetry matrix and another proportional to an skew symmetry one. In this way, the residual part equals a rotation of a gradient of a function, i.e., $\nabla\mathbf{V}$, which is given by $\vec{v}=\kappa\Omega\nabla\mathbf{V}$, $\Omega$ is the unit skew-symmetry, $\kappa$ a set of parameters, which guaranties the orthogonality condition is fulfilled. Accordingly, Graham obtains[8] (see Equ.(4)):
[…]
$\omega\wedge d\omega=(\sum^{N}_{l=1}f_{l}dx_{l})\wedge(\sum^{N}_{k,i=1}\frac{\partial f_{i}}{\partial x_{k}}dx_{k}dx_{i})=\sum^{N}_{l,k,i=1}f_{l}\frac{\partial f_{i}}{\partial x_{k}}dx_{l}dx_{k}dx_{i})=0,$ (35)
which means that the vector function $\vec{f}$ is orthogonal to its $rotor"$. In fact, previous relation should be checked before getting involved in the Graham scheme. In all the examples where the scheme was successfully applied [9], previous condition must have been satisfied. In other words, the Graham scheme applies for the special class of functions where the p=1 diff-form defined in Eqn.(14) satisfies Eqn.(35).
VII Acknowledgments.
We thank Prof. Dr. Thomas Klinker, Universitaet Hamburg, Germany, for donating the important book for this work on Differential Forms by Henry Cartan and for financial support for this project. We thank Dr. Hector Giusti, DIC USB, for donating a PC which was used exclusively for writing this work.